OpenStudy (anonymous):

1. Write the equation of the line through the points (1, -2) and (4, 1) 2. Write the equation of the arithmetic sequence that starts with -7 and each term decreases by 3 a) y = 1/2x +1 b) y - 3 = 1/2 (x -4) c) x - 2y = -2 d) f(0) = 1, f(n) = f(n - 1) + 1/2 1. Verify that the four equations above are equivalent. 2. Explain how you know that the four equations are linear. 3. You have been appointed as a mathematics efficiency expert. Your job is to compare these four forms of equations for different uses and decide which form is most efficient and effective for each use. CONTINUED ↓ ↓

OpenStudy (anonymous):

The investigation will be conducted in four parts with a report to be written at the end. 4. What is the important difference between the type of situations that can be modeled with a recursion formula and the other equation forms?

OpenStudy (anonymous):

Sorry I just moved to a new school in the middle of the semester, so I don't understand any of this :(

OpenStudy (anonymous):

for the first one: first find the slope. y2-y1/x2-x1 = 1-(-2)/ 4-1= 4/3 then you must plug in the points and the slope into y-y1=m(x-x1) y-1=3/4(x-(-2)) y-1=3/4x- 6/4 y-1=3/4x-3/2 add 1 on both sides y=3/4x-3/2+1 since you cannot add fractions that don't have a common denominator, you must multiply the top and the bottom by 2 3/2 + 2/2 = 5/2 so therefore the equation is: y=3/4x+ 5/2 I believe that is right.

OpenStudy (anonymous):

Thank you sooo much @pinkanchor212

OpenStudy (anonymous):

no problem!